You are given the following cost​ data: Total fixed costs are $60.
Answer #1

In a competitive market, the firm will produce that output where price equals the MC. Hence, the firm will produce 5units of output because this is where price equals Marginal Cost. For producing 5 units of output, the firm will incur a variable cost of 255 units. For producing 4units the variable cost will be 165 units. Hence, the marginal cost will be 255-165 = 90 units. Fixed cost is a fixed number, hence it's MC will be zero. As the price of the output is 90 and the MC is also 90 for 5 units. Hence, 5 units will be produced.

Total Revenue = Price * Quantity = 90*5 = 450 units.

Total Cost = Fixed Cost + Variable Cost = 60+255= 315 units.

Profit = Total Revenue - Total Cost = 450 - 315 = 135 units.

In the short run, the firm will have two two equilibrium conditions : (a) MC = MR , and (b) MC curve must be rising at the point of equilibrium.

In the long run, the firm will increase its output by changing their capital equipment by expanding the output or by replacing the old lower capacity plants. In the long run for the firm to be in equilibrium, besides marginal cost being equal to price, the price must be equal to average cost.